OpenStudy (anonymous):

Trying to study for a unit test tomorrow, refresh me on this please!! Write an equation of the line, in point-slope form, that passes through the two given points. points: (–2, 15), (9, –18)

Directrix (directrix):

Do you know how to compute the slope of the line determined by the two points?

OpenStudy (anonymous):

Is it the m=y1-x1/y2-x2

Directrix (directrix):

Almost. See the attached file. Think of slope as change in y over change in x. So, m = (y1 - y2) / (x1 - x2) where (x1,y1) and (x2, y2) are points on the line.

OpenStudy (anonymous):

So will it look like: m=(-2-9)/(15-(-18))

Directrix (directrix):

So, slope here is m = [(15) - (-18) / [(-2) - (9)]. How does yours differ from mine. I'm not saying mine is for sure correct.

Directrix (directrix):

The difference in the y values goes on top. Think of slope as change in y over change in x.

OpenStudy (anonymous):

Looks similar just more parenthesis, and organized differently

OpenStudy (anonymous):

y-2=-1/3(x+15) is that right?!?!

Directrix (directrix):

Your slope = -1/3. My slope = -3. You flipped the formula. It is the difference in the y values divided by the difference in the x values. You have the difference in the x values divided by the difference in the y values. That is not the correct formula. It is not the correct slope.

Directrix (directrix):

Do you see that?

Directrix (directrix):

Now, use y - y1 = x - x1 to get the equation of the line. Let (–2, 15) be (x1, y1). @kim2012 Question?

Directrix (directrix):

I'm waiting to hear from you.

OpenStudy (anonymous):

y-2=-3(x-15)

Directrix (directrix):

What does that simplify to? y-2=-3(x-15) Or, is this point/slope form? I was thinking slope/intercept form but that is not what was asked.

OpenStudy (anonymous):

Im trying to find out the eqaution i would use in order to find those two points. So if the slope it 3 then it think its that or y-15=-3(x+2)

Directrix (directrix):

Now, use y - y1 = x - x1 to get the equation of the line. Let (–2, 15) be (x1, y1). ============ y - (-15) = -3(x - (-2) ) y + 15 = -3(x + 2) y + 15 = -3x - 6----> ========== y-2=-3(x-15) Do you see where your sign error happened?

Directrix (directrix):

Okay, I hear you. There are several ways to write the same equation. But, you have sign errors to see about and also you need to practice the slope formula.

Directrix (directrix):

One nice thing about Algebra is that it is not a guessing game. So, if you use the formula correctly, you should get the equation: y - (-15) = -3(x - (-2) ) y + 15 = -3(x + 2) y + 15 = -3x - 6---->

Directrix (directrix):

Show the steps. These negative terms when subtracted can be tricky.